EXPERIMENTAL STUDY OF THE Cs_{2} TRIPLET STATES BY PERTURBATION FACILITATED INFRARED-INFRARED DOUBLE RESONANCE AND TWO-PHOTON EXCITATION SPECTROSCOPY

\underline{FENG XIE}, DAN LI, LI LI, Department of Physics and Key Laboratory of Atomic and Molecular Nanosciences, Tsinghua University,Beijing 100084, China; VLADIMIR B. SOVKOV, VALERY S. IVANOV, V. A. Fock Institute of Physics, St. Petersburg State University, 1 Ulyanovskaya Street, Petrodvorets, St. Petersburg 198504, Russia; and SYLVIE MAGNIER, Laboratoire de Physique des Atomes, Lasers, Molećule et Surfaces (PALMS), CNRS et Universite´ RENNES I (UMR 6627), Campus de Beaulieu, Bt 11B 35042 Cedex, France.

Fifty 3^{3}Σ_{g}^{+} v= 2 − 9 ← X^{1}Σ_{g}^{+} two-photon transitions of Cs_{2} have been observed in the infrared wavelength. Vibrational resolved fluorescence into both bound levels and the continuum of the a^{3}Σ_{u}^{+} state has been recorded with monochromator. Calculated 3^{3}Σ_{g}^{+} → a^{3}Σ_{u}^{+} fluorescence spectra are compared with observed spectra. With two infrared diode lasers,totally 43 ro-vibrational levels of the Cs_{2} 3^{3}Π_{g} state have been observed by perturbation facilitated infrared-infrared double resonance spectroscopy. Rotationally resolved fluorescence to the a^{3}Σ_{u}^{+} state was also recorded.